10–15 IVPs Solve the following initial value problems. 10. yı = -4yı + 5y2 yż = -yı + 2y2 Yı(0) = 0, y2(0) = 4
Answer #1

. 10 fet A=1 1-4 57 The characteristie polynomical of Ais A(1) = 72-lfaceA)+ delt =4*+21-3 A1A)=0 => 72+27-3=0 => 6+3)(2-1) =5a45b=0 fel b=1, then a=!! ! : ) General solution of the given System is your soul - gent as ek nie -34 t7 5ge the 34 it TGE